Suppose the market for watermelon can be described as the graph below . If Jon is willing to pay 8 dollars


1
Expert's answer
2022-03-03T07:24:46-0500

Consumer surplus= "\frac{1}{2}\times Q \times (P_2-P_2)"

"\frac{1}{2}\times 6\times (8-5)" = 9